Grade 12Differential Calculus

Q1 FIND OUT FOR WHAT INTEGRAL VALUES OF N THE NUMBER VALUES OF N THE NUMBER 3 IS A PERIOD OF THE FUNCTION ; f(x)= COSNX.SIN(5/X)

Q2 FIND THE INVERSE OF f(x) = 2LOGX+8. &HENCE SOLVE THE EQUATION f(x)= f -1(x)

tapish khari
16 Years agoGrade 12

1 Answer

mounika gajawada
15 Years ago

i'm unable to understand ur first ques.......plz post it properly

 

ans for Q2 is x=10

let 2logx+8=y

→ logx=log(y-8)/log2

→ x=10{log(y-8)/log2}

→ f-1(x) = 10{log(y-8)/log2}

f(x)=f-1(x) only wen y=x

→ x= 2logx +8 

→ x=10